OpenStudy (anonymous):

find a quadratic equation with the following roots . 5 +2i and 5-2i

OpenStudy (anonymous):

\[x^2-10 x+29\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

when we say a quadratic equation has roots a and b, we mean that the values a and b "satisfy" the equation. in other words, if you substitute x with either a or b, the resulting value of the function is 0. this is why roots are also called zeros

OpenStudy (anonymous):

so if a quadratic equation has roots a and b, the quadratic equation is (x-a)(x-b)=0. notice that if you substitute x with a, the expression becomes (a-a)(a-b) which is equal to 0. similarly when you substitute b for x you get (b-a)(b-b) which is also equal to 0

OpenStudy (anonymous):

(x-a)(x-b) = 0 can be expanded as x*x -a*x-b*x +a*b =0 x^2 -(a+b)x +ab = 0 so if your roots are 5+2i and 5-2i, the resulting quadratic equation is x^2 - (5+2i + 5-2i)x +(5+2i)(5-2i) = 0 which is x^2 - (10)x + (5^2-(2i)^2) = 0 x^2 -10x +(25 - (-4)) =0 x^2 -10x +29 = 0